OpenStudy (anonymous):

What is the rate of change for the linear relationship modeled in the table? x y −1 10 1 9 3 8 5 7

Mehek (mehek14):

\(\dfrac{7 - 8}{5 - 3}=\dfrac{-1}{2}\)

OpenStudy (jadedinsomniac):

SWIGGITY SWOOGITY COMMIN FOR THAT BOOTY HEAVY

OpenStudy (anonymous):

precisely

OpenStudy (jadedinsomniac):